OpenStudy (katielong):

"Jack spends 1 1/2 as long on his homework as Jill. Last week Jill spent 9 1/2 hours doing homework. How long did Jack spend doing homework?"

OpenStudy (pfenn1):

If x = time Jack spends on homework y=time Jill spends on homework x=(1 1/2) * y y= 9 1/2 hours solve for x

OpenStudy (katielong):

ahh so its jacks times Jills so i would have to divide Jills by 1 1/2? :)

OpenStudy (pfenn1):

If Jack spends 1 1/2 as much time as Jill, you would multiply Jill's time by 1 1/2. So if Jill spends an hour doing homework, Jack spends 1 1/2 hours. So if Jill spends 9 1/2 hours doing homework, how much time does Jack spend?

OpenStudy (katielong):

1/2 longer than Jill

OpenStudy (pfenn1):

yes, how many hours does Jack spend if Jill spends 9 1/2 hours?

OpenStudy (katielong):

well the question initially said 1 1/5 not 1 1/2 so the answer to that would be 7 11/12??

OpenStudy (pfenn1):

If Jack spends 1 1/5 as much time as Jill and Jill spends 9 1/2 hours: Jack's time = (1 1/5) times (9 1/2) = ? The answer must be greater than 9 1/2 hours.

OpenStudy (katielong):

ahhhh :S whats the answer ?

OpenStudy (katielong):

11.4

OpenStudy (pfenn1):

Yes! Correct!

OpenStudy (katielong):

so 11 2/5

OpenStudy (pfenn1):

Yep

OpenStudy (katielong):

part b = "the week before, jack spent 21 3/4 hours on homework. how long did Jill spend?"

OpenStudy (pfenn1):

Assuming Jack still takes 1 1/5 times as long as Jill, you can use the same equation as the last time x= (1 1/5) * y where x is Jack's time and y is Jill's time y = x/ (1 1/5) = (21 3/4) / ( 1 1/5) = ? you find Jill's time by dividing Jack's time by 1 1/5.

OpenStudy (katielong):

okay i'll tryy :/

OpenStudy (pfenn1):

You can do it.

OpenStudy (katielong):

i cant!!!!

OpenStudy (katielong):

843/4 / 6/5 then 843/4 x 5/6 am i getting there?

OpenStudy (pfenn1):

very close....21 3/4 = 87/4 and then you do as you wrote 87/4 x 5/6 = answer!

OpenStudy (katielong):

ahhh, slipped up slightly :L the answer is... 75/24 = 3 3/24 = 3 1/8

OpenStudy (katielong):

or 21 3/4 - 3 1/8

OpenStudy (pfenn1):

I got 87/4 x 5/6 = 435/24 = 18 3/4

OpenStudy (katielong):

ah right, i see where i went wrong... Thanks :D Id give you 100 medals if i could !!! :)

OpenStudy (pfenn1):

Thanks but not necessary.

OpenStudy (katielong):

whattt!!! it was wrong!! (the we we just did )

OpenStudy (pfenn1):

Shoot. I will check my math again.

OpenStudy (katielong):

yeh but that means i need to do all of it again becasue its on mymaths :SSSSSSS!!!!

OpenStudy (pfenn1):

I just did it another way and got 17 2/5. Does that seem right?

OpenStudy (katielong):

im not sure, im doing it again now but its a different question this time :L

OpenStudy (katielong):

whats 16.1388888 reccurring as a fraction?

OpenStudy (pfenn1):

I checked up my math and this time I got 18 1/8.

OpenStudy (katielong):

doesn't matter now though!!

OpenStudy (pfenn1):

I can't find an answer to the recurring decimal to fraction question.

OpenStudy (katielong):

its fine